### v2

@@ Abstract
With the proposal of the Segment Anything Model (SAM), fine-tuning SAM for medical image segmentation (MIS) has become popular. However, due to the large size of the SAM model and the significant domain gap between natural and medical images, fine-tuning-based strategies are costly with potential risk of instability, feature damage and catastrophic forgetting. Furthermore, some methods of transferring SAM to a domain-specific MIS through fine-tuning strategies disable the model’s prompting capability, severely limiting its utilization scenarios. In this paper, we propose an Auto-Prompting Module (APM), which provides SAM-based foundation model with Euclidean adaptive prompts in the target domain. Our experiments demonstrate that such adaptive prompts significantly improve SAM’s non-fine-tuned performance in MIS. In addition, we propose a novel non-invasive method called Incremental Pattern Shifting (IPS) to adapt SAM to specific medical domains. Experimental results show that the IPS enables SAM to achieve state-of-the-art or competitive performance in MIS without the need for fine-tuning. By coupling these two methods, we propose ProMISe, an end-to-end non-fine-tuned framework for Promptable Medical Image Segmentation. Our experiments demonstrate that both using our methods individually or in combination achieves satisfactory performance in low-cost pattern shifting, with all of SAM’s parameters frozen.
Keywords:
SAM Medical image segmentation Pattern shifting
✉  The corresponding authors*   These authors contributed equally to this workThe code can be found at https://github.com/xinkunwang111/ProMISe
Figure 1: Overview of ProMISe. All three components of the original SAM are frozen. The Auto-Prompting Module (APM) leverages features from the image encoder to predict optimal prompts in Euclidean space. The APM can be implemented using various operators and modules, such as CNN and Transformer. The Pattern Embedding (PaE) module analyzes the image embedding to extract pattern gaps between the target and source domains. The Incremental Pattern Shifting (IPS) tokens are added to the mask tokens of the output tokens to realize the decoder’s shifting of mask patterns.
@@ 1 Introduction
As demonstrated in Fig. 1, based on the above arguments: 1) We propose the Auto-Prompting Module (APM), which leverages the SAM framework for training and provides adaptive Euclidean prompts for SAM; 2) We propose Incremental Pattern Shifting (IPS), a novel non-invasive pattern shifting method which couples a Pattern Embedding (PaE) module with IPS tokens to cost-effectively shift the prior pattern knowledge of the mask decoder. The IPS method enables SAM to achieve state-of-the-art (SOTA) and competitive performance without the need for fine-tuning; 3) We couple the above two methods to propose the Promptable Medical Image Segmentation (ProMISe) framework utilizing SAM. Notably, we are the first to propose methods for transferring SAM to MIS while keeping all of SAM’s parameters frozen, resulting in significantly reduced training costs, improved stability, and the ability to retain spatial prompting capability.
2 Method
2.1 Adaptive Prompt
@@
As shown in Fig. 1, we design a lightweight end-to-end module, called Auto-Prompting Module (APM), which effectively integrates the multi-level features of SAM’s image encoder to predict optimal prompts in Euclidean space. By generating adaptive prompts that provide more fine-grained spatial information, the APM significantly enhances the untrained performance of SAM in the target domain. This indicates the highly expressive feature extraction capability of SAM, even in previously unseen domains.
@@
2.2 Incremental Pattern Shifting (IPS)
@@
After rethinking the components of the SAM, we observe that the SAM has a very efficient mask decoder that refers to DETR [14] and MaskFormer [15]. In this study, we find that in the mask decoder, output tokens initially receive the location information of the point of interest (PoI) or region of interest (RoI) from the prompt encoder through self-attention. Subsequently, as shown in Fig. 1, output tokens extract semantic information from the image embedding through cross-attention, forming semantic patterns. Based on these findings, we argue that with a large amount of training, the output tokens can be regarded as the pattern information acquired from the source domain, enabling SAM to generate predicted masks based on PoI or RoI. Therefore, assuming the image encoder of SAM is sufficiently powerful and well-trained, transferring SAM to MIS can be equivalent to transferring the mask prediction patterns of SAM to MIS.
@@
With the above arguments, we propose a method called Incremental Pattern Shifting (IPS). As shown in Fig. 1, we extract the pattern shifting information (IPS tokens) from the image embedding using a lightweight PaE module, and then non-invasively shift the patterns of the mask decoder by adding the IPS tokens to the mask tokens (Eqs. 1-3). In Fig. 2, the mask decoder, when provided with the same PoI, generates mask predictions that are more compatible with the target domain after pattern shifting. Notably, our experiments show the effectiveness of IPS, utilizing GT-based prompts for both training and inference, in enabling SAM to achieve SOTA or competitive performance in MIS tasks.
2.3 ProMISe framework
Our proposed IPS method effectively facilitates segmentation pattern shifting. However, using GT-based prompts for pattern transfer does not allow end-to-end training. In addition, GT-based prompt sampling introduces significant randomness, thereby posing challenges to the stability of the training process. Therefore, we couple the IPS method with APM to achieve end-to-end pattern shifting for SAM to target medical domains. Moreover, favored by the preservation of Euclidean spatial form prompts, the ProMISe framework can handle both automatic and manual prompts during inference while retaining its interpretability.
@@ 3.1 Experimental Setup
We individually conduct experiments in two modalities: endoscopy and dermoscopy. For endoscopy experiments, we follow the experimental setups of PraNet [16], DuAT [24], and SSFormer [17]. We extract 1450 images from Kavsir [18] and CVC-ClinicDB [21] as the training set. Tests are then conducted on EndoScene [22], Kvasir, CVC-ColonDB [19], and ETIS [23], using Mean Absolute Error (MAE), mean Dice, and mean IoU as metrics. For dermatoscopy, we use ISIC2018 for training and testing, with mean Dice and mean IoU as metrics. We utilize ViT-B as the image encoder for all SAM-related methods in the experiments. In addition, we apply three prompt settings in training and inference, namely 3/5/16 points (1/2/8 positive + 2/3/8 negative, 3/5/16P, respectively).
We implement our methods in pytorch, using one TESLA A100 GPU. Adam optimizer and a learning rate of 0.00001 are employed. The training period is 200 epochs. Our loss is a combination of Dice and BCE loss.
Table 1: Quantitative comparison of the performance enhancement brought to SAM by adaptive prompts provided by APMs (RN and Cross). ${}^{a}$ Trainable Parameters.
@@
[…]
@@
3.2 Adaptive Prompt
As described in Sect. 2.1, to study the actual performance of vanilla SAM in MIS, we train and test two types of APMs, i.e., ResNet34 (RN) and a one-layer cross-attention Transformer module (Cross), on polyp benchmarks. As shown in Tab. 1, the adaptive prompts generated by both APMs effectively improve the performance of SAM in the polyp segmentation task, achieving a level comparable to the baseline MIS segmentation model. Furthermore, we observe a significant reduction in MAE by increasing the number of point prompts with more provided fine-grained information, which suggests that our method effectively utilizes the boundary sensitivity of vanilla SAM.
Table 2: Quantitative comparison of our IPS, SOTA methods, and other SAM-based methods. Best-in-class results are bolded, while second-best results are underlined. The 3P|5P and 3P|16P represent training our IPS on 3 GT-based prompt points and testing them with 5 and 16 ones. ${}^{a}$ Number of Trainable Parameters. ${}^{b}$ Number of Endoscopy/Dermoscopy images is utilized in the corresponding training set. $*{}^{*}$ Some test set images from the corresponding dataset may be seen in the training process.
Benchmarks Kvasir EndoScene ColonDB ETIS ISIC2018
[…]
@@ 3.2 Adaptive Prompt
IPS-16P 1.3M 1450/2594 0.911 0.851 0.897 0.821 0.874 0.789 0.840 0.750 0.921 0.850
Figure 3: Qualitative comparison of the performance with same prompts in ColonDB.
3.3 Pattern Shifting
In Tab. 2, we train the proposed IPS method (PaE with 4 IPS tokens) using GT-based prompt points. We perform tests using the same GT-based point prompts for SAM-related models and compare our method (SAM with IPS) to SOTA methods, SAM and SAM-Med2D. As shown in Tab. 2, our proposed IPS significantly improves the performance of SAM on benchmarks, achieving mDice improvements of 10%-25% in Kvasir, 19%-31% in EndoScene, 25%-33% in ColonDB, 25%-32% in ETIS, and 18%-37% in ISIC2018. Compared to SAM-Med2D, which is trained on a large-scale medical dataset and utilizes a complex training strategy, our method also achieves a remarkable improvement. Experimental results indicate that the IPS method adequately leverages SAM’s powerful generalization and feature expression capabilities. Even in the unfamiliar and challenging benchmarks of ColonDB and ETIS, IPS enables SAM to achieve SOTA performance. Using SSFormer as a baseline, we obtain up to 10% and 7% improvement in ColonDB and ETIS, respectively. It is promising that the above performance improvement depends only on a tiny number of trainable parameters (1.3M), demonstrating that our IPS method is very lightweight and efficient.
Table 3: Performance of SAM with ProMISe (Cross and RN) is tested using GT-base prompts. ${}^{a}$ Number of Trainable Parameters. ${}^{b}$ Number of Endoscopy/Dermoscopy images is utilized in the corresponding training set. $*{}^{*}$ Some test set images from the corresponding dataset may be seen in the training process.
Benchmarks Kvasir EndoScene ColonDB ETIS ISIC2018
@@ 3.3 Pattern Shifting
U-Net - 1450/2594 0.818 0.746 0.710 0.627 0.512 0.444 0.398 0.335 0.855 0.785
ResUNet++ - 1450/2594 0.821 0.743 0.707 0.624 0.483 0.410 0.401 0.344 0.809 0.729
SAM-5P - - 0.750 0.645 0.656 0.582 0.569 0.482 0.541 0.472 0.687 0.569
SAM-16P - - 0.719 0.620 0.692 0.613 0.548 0.467 0.524 0.455 0.738 0.624
Med2D-5P 184.5M 5838/7935 $*{}^{*}$0.822 $*{}^{*}$0.735 0.722 0.623 0.686 0.576 0.677 0.571 $*{}^{*}$0.884 $*{}^{*}$0.813
Med2D-16P 184.5M 5838/7935 $*{}^{*}$0.832 $*{}^{*}$0.748 0.727 0.620 0.685 0.575 0.622 0.514 $*{}^{*}$0.893 $*{}^{*}$0.823
Cross-5P 45.6M 1450/2594 0.834 0.744 0.788 0.687 0.744 0.636 0.678 0.578 0.742 0.631
Cross-16P 45.6M 1450/2594 0.858 0.777 0.768 0.661 0.732 0.626 0.691 0.591 0.819 0.705
RN-5P 23.0M 1450/2594 0.776 0.673 0.705 0.587 0.664 0.547 0.605 0.508 0.798 0.688
RN-16P 23.0M 1450/2594 0.846 0.759 0.803 0.735 0.735 0.624 0.673 0.566 0.878 0.791
As demonstrated in Fig. 3 and Tab. 2, SAM and SAM-Med2D struggle to handle exposures and insignificant small-volume targets well, even when provided with more fine-grained guidance through an increased number of prompt points. In contrast, SAM with IPS significantly optimizes its predicted masks through non-invasive pattern shifting based on the same prompts (Fig. 3).
@@ 3.4 ProMISe framework
To avoid random sampling of prompts and thus achieve end-to-end pattern shifting, we propose the ProMISe framework, which couples the APM and IPS for training with adaptive prompts. When tested using GT-based prompts, the end-to-end pattern shifting of ProMISe with both APMs (Cross and RN) significantly improves SAM’s performance to a practical and competitive level in MIS (Tab. 3). Importantly, ProMISe maintains its interpretability using both adaptive/GT-based Euclidean prompts and keeps all of SAM’s parameters frozen, resulting in a more practical and applicable approach for real clinical scenarios.

### v3

@@ Abstract
With the proposal of Segment Anything Model (SAM), fine-tuning SAM for medical image segmentation (MIS) has become popular. However, due to the large size of the SAM model and the significant domain gap between natural and medical images, fine-tuning-based strategies are costly with potential risk of instability, feature damage and catastrophic forgetting. Furthermore, some methods of transferring SAM to a domain-specific MIS through fine-tuning strategies disable the model’s prompting capability, severely limiting its utilization scenarios. In this paper, we propose an Auto-Prompting Module (APM), which provides SAM-based foundation model with Euclidean adaptive prompts in the target domain. Our experiments demonstrate that such adaptive prompts significantly improve SAM’s non-fine-tuned performance in MIS. In addition, we propose a novel non-invasive method called Incremental Pattern Shifting (IPS) to adapt SAM to specific medical domains. Experimental results show that the IPS enables SAM to achieve state-of-the-art or competitive performance in MIS without the need for fine-tuning. By coupling these two methods, we propose ProMISe, an end-to-end non-fine-tuned framework for Promptable Medical Image Segmentation. Our experiments demonstrate that both using our methods individually or in combination achieves satisfactory performance in low-cost pattern shifting, with all of SAM’s parameters frozen. Code is available at https://github.com/xinkunwang111/ProMISe
@@ 1 Introduction
Figure 1: Overview of ProMISe. All three components of the original SAM are frozen. The Auto-Prompting Module (APM) leverages features from the image encoder to predict optimal prompts in Euclidean space. The Pattern Embedding (PaE) module analyzes the image embedding to extract pattern gaps between the target and source domains. The Incremental Pattern Shifting (IPS) tokens are added to the mask tokens of the output tokens to realize the decoder’s shifting of mask patterns.
Figure 2: Detailed structure of APM and PaE. The APM can be implemented using various operators and modules, such as CNN and Transformer.
@@
As demonstrated in Fig. 1 and Fig. 2, based on the above arguments: 1) We propose the Auto-Prompting Module (APM), which leverages the SAM framework for training and provides adaptive Euclidean prompts for SAM; 2) We propose Incremental Pattern Shifting (IPS), a novel non-invasive pattern shifting method which couples a Pattern Embedding (PaE) module with IPS tokens to cost-effectively shift the prior pattern knowledge of the mask decoder. The IPS method enables SAM to achieve state-of-the-art (SOTA) and competitive performance without the need for fine-tuning; 3) We couple the above two methods to propose the Promptable Medical Image Segmentation (ProMISe) framework utilizing SAM. Notably, this method is able to transfer SAM to MIS while keeping all of SAM’s parameters frozen, resulting in significantly reduced training costs, improved stability, and the ability to retain spatial prompting capability.
2 Adaptive Prompt
2.1 Motivation
@@
2.2 Proposed Method
As shown in Fig. 1 and 2, we design a lightweight end-to-end module, called Auto-Prompting Module (APM), which effectively integrates the multi-level features of SAM’s image encoder to predict optimal prompts in Euclidean space. By generating adaptive prompts that provide more fine-grained spatial information, the APM significantly enhances the untrained performance of SAM in the target domain. This indicates the highly expressive feature extraction capability of SAM, even in previously unseen domains.
@@ 2.2 Proposed Method
3 Incremental Pattern Shifting
3.1 Rethink the mask decoder
@@
After rethinking the components of the SAM, we observe that the SAM has a very efficient mask decoder that refers to DETR [2] and MaskFormer [5]. In this study, we find that in the mask decoder, output tokens initially receive the location information of the point of interest (PoI) or region of interest (RoI) from the prompt encoder through self-attention. Subsequently, as shown in Fig. 1, output tokens extract semantic information from the image embedding through cross-attention, forming semantic patterns.
3.2 Proposed Method
Based on these findings, we argue that with a large amount of training, the output tokens can be regarded as the pattern information acquired from the source domain, enabling SAM to generate predicted masks based on PoI or RoI. Therefore, assuming the image encoder of SAM is sufficiently powerful and well-trained, transferring SAM to MIS can be equivalent to transferring the mask prediction patterns of SAM to MIS.
@@ 3.2 Proposed Method
With the above arguments, we propose a method called Incremental Pattern Shifting (IPS). As shown in Fig. 1, we extract the pattern shifting information (IPS tokens) from the image embedding using a lightweight PaE module, and then non-invasively shift the patterns of the mask decoder by adding the IPS tokens to the mask tokens (Eqs. 1-3). In Fig. 3, the mask decoder, when provided with the same PoI, generates mask predictions that are more compatible with the target domain after pattern shifting. As an experimental validation of Fig. 3, the right two columns of Fig. 4 (SAM and IPS) show that the transfer of the segmentation pattern of SAM using only IPS is solid. Notably, our experiments show the effectiveness of IPS, utilizing GT-based prompts for both training and inference, in enabling SAM to achieve SOTA or competitive performance in MIS tasks.
4 ProMISe framework
As shown in Fig. 1, IPS does not conflict with the proposed APM which provides prompts for end-to-end SAM transfer and inference. Deploying non-invasive APM, PaE, and IPS tokens to SAM allows automation and transfer of original SAM to MIS tasks. Our proposed IPS method effectively facilitates segmentation pattern shifting. However, using GT-based prompts for pattern transfer does not allow end-to-end training process. In addition, GT-based prompt sampling may introduce randomness, thereby posing challenges to the stability of the training and evaluation process. Therefore, we couple the IPS method with APM to achieve end-to-end pattern shifting for SAM to target medical domains. Moreover, favored by the preservation of Euclidean spatial form prompts, this ProMISe framework can handle both automatic and manual prompts during inference while retaining its interpretability.
@@ 5.1 Experimental Setup
We individually conduct experiments in two modalities: endoscopy and dermoscopy. For endoscopy experiments, we follow the experimental setups of PraNet [8], DuAT [21], and SSFormer [23]. We extract 1450 images from Kavsir [11] and CVC-ClinicDB [19] as the training set. Tests are then conducted on Kvasir, EndoScene [20], CVC-ColonDB [1], and ETIS [22], using Mean Absolute Error (MAE), mean Dice, and mean IoU as metrics. For dermatoscopy, we use ISIC2018 for training and testing, with mean Dice and mean IoU as metrics. We utilize ViT-B as the image encoder for all SAM-related methods in the experiments. In addition, we apply three prompt settings in training and inference, namely 3/5/16 points (1/2/8 positive + 2/3/8 negative, 3/5/16P, respectively). All GT-based prompt points used in this paper were obtained by random sampling from GT masks. For fair comparisons, the prompt points provided to the models are exactly the same in each experimental setting.
We implement our methods in pytorch, using one TESLA A100 (80G) GPU. Adam optimizer and a learning rate of 0.00001 are employed. The training period is 200 epochs. Our loss is a combination of Dice and BCE loss. The IPS method requires approximately 22 GPU hours and 25 GPU hours for training in polyps and ISIC, respectively. In addition, APM-cross requires approximately 28 GPU hours and 30 GPU hours for training in polyps and ISIC. The training periods for APM-RN are 21 GPU hours and 25 GPU hours.
Figure 4: Comparison of the performance with different prompt number in endoscopy datasets. The 3P, 5P and 16P contain 1, 2, and 8 positive, as well as 2, 3, and 8 negative points, respectively.
5.2 Adaptive Prompt
As described in Sect. 2, to study the actual performance of vanilla SAM in MIS, we train and test two types of APMs, i.e., ResNet34 (RN) and a one-layer cross-attention Transformer module (Cross), on polyp benchmarks. As shown in Tab. 1, the adaptive prompts generated by both APMs effectively improve the performance of SAM in the polyp segmentation task, achieving a level comparable to the baseline MIS segmentation model. Furthermore, we observe a significant reduction in MAE by increasing the number of point prompts with more provided fine-grained information, which suggests that our method effectively utilizes the boundary sensitivity of vanilla SAM.
Table 1: Comparison of the performance enhancement brought to SAM by adaptive prompts provided by APMs (RN and Cross). a Trainable Parameters.
@@ 5.2 Adaptive Prompt
Methods TPa mDice mIoU MAE↓ mDice mIoU MAE↓ mDice mIoU MAE↓
@@
5.3 Pattern Shifting
In Tab. 2 and 3, we train the proposed IPS method (PaE with 4 IPS tokens) using GT-based prompt points. We perform tests using the same GT-based point prompts for SAM-related models and compare our method (SAM with IPS) to SOTA methods, SAM and SAM-Med2D. Following the mainstream validation frameworks in the field of polyp segmentation [8, 23]. the Kvasir and EndoScene benchmarks are used to validate the transfer performance of the proposed method, while the ColonDB and ETIS benchmarks are used to validate the generalization ability of the proposed method due to the domain gap with the training set.
Table 2: Quantitative comparison of our IPS, SOTA methods, and other SAM-based methods. Best-in-class results are bolded, while second-best results are underlined. The 3P|5P and 3P|16P represent training our IPS on 3 GT-based prompt points and testing them with 5 and 16 ones. a Number of Trainable Parameters. b Number of Endoscopy/Dermoscopy images is utilized in the corresponding training set. Some test set images from the corresponding dataset may be seen in the training process.
Benchmarks Kvasir EndoScene ISIC2018
Methods TPa Imagesb mDice mIoU mDice mIoU mDice mIoU
@@ 5.3 Pattern Shifting
IPS-3P 1.3M 1450/2594 0.797 0.704 0.806 0.718 0.854 0.762
IPS-3P|5P 1.3M 1450/2594 0.821 0.732 0.804 0.716 0.865 0.774
IPS-3P|16P 1.3M 1450/2594 0.843 0.752 0.815 0.721 0.874 0.785
@@
IPS-16P 1.3M 1450/2594 0.902 0.835 0.888 0.810 0.915 0.847
As shown in Tab. 2, our proposed IPS significantly improves the performance of SAM on familiar benchmarks, achieving mDice improvements of 10%-21% in Kvasir, 19%-29% in EndoScene, and 18%-37% in ISIC2018. Compared to SAM-Med2D, which is trained on a large-scale medical dataset and utilizes a complex training strategy, our method also achieves a remarkable improvement.
Experimental results in Tab. 3 indicate that the IPS method adequately leverages SAM’s powerful generalization and feature expression capabilities. Even in the unfamiliar and challenging benchmarks of ColonDB and ETIS, IPS enables SAM to achieve SOTA performance. Using SSFormer as a baseline, we obtain up to 10% and 9% improvement in ColonDB and ETIS, respectively. Moreover, compared to vanilla SAM, IPS can achieving mDice improvements of 25%-33% in ColonDB, 20%-33% in ETIS. It is promising that the above performance improvement depends only on a tiny number of trainable parameters (1.3M). Those demonstrate that the IPS can efficiently transfer SAM to the target task and achieve SOTA-level performance with powerful generalization ability. This makes the low-cost rapid deployment of SAM promising.
Table 3: Generalization performance comparison of our IPS, SOTA methods, and other SAM-based methods. Best-in-class results are bolded, while second-best results are underlined.
Benchmarks ColonDB ETIS
Methods TPa Imagesb mDice mIoU mDice mIoU
SOTA Methods
U-net [17] - 1450/2594 0.512 0.444 0.398 0.335
PraNet [8] - 1450/2594 0.712 0.640 0.628 0.567
[…]
SAM-based Methods
SAM-3P [12] - - 0.447 0.356 0.464 0.381
SAM-5P - - 0.569 0.482 0.541 0.472
SAM-16P - - 0.548 0.467 0.524 0.455
Med2D-3P [7] 184.5M 5838/7935 0.689 0.588 0.633 0.524
Med2D-5P 184.5M 5838/7935 0.686 0.576 0.677 0.571
Med2D-16P 184.5M 5838/7935 0.685 0.575 0.622 0.514
IPS-3P 1.3M 1450/2594 0.724 0.618 0.659 0.569
IPS-3P|5P 1.3M 1450/2594 0.761 0.655 0.727 0.626
IPS-3P|16P 1.3M 1450/2594 0.783 0.678 0.763 0.674
[…]
IPS-16P 1.3M 1450/2594 0.874 0.789 0.854 0.770
As demonstrated in Fig. 4 and Tab. 2 and 3, SAM and SAM-Med2D struggle to handle exposures and insignificant small-volume targets well, even when provided with more fine-grained guidance through an increased number of points (i.e. more prompt points). In contrast, our proposed IPS significantly optimizes its predicted masks through non-invasive pattern shifting based on the same prompts (Fig. 4). Furthermore, the results in Tab. 2 and 3 demonstrate that even if the number of prompt points used in training does not match the number of points used in testing, the IPS can still handle these information gaps well.
@@ 5.4 ProMISe framework
To avoid random sampling of prompts and thus achieve end-to-end pattern shifting, we propose the ProMISe framework, which couples the APM and IPS for training with adaptive prompts. When tested using GT-based prompts, the end-to-end pattern shifting of ProMISe with both APMs (Cross and RN) significantly improves SAM’s performance to a practical and competitive level in MIS (Tab. 4 and 5). Importantly, ProMISe maintains its interpretability using both adaptive/GT-based Euclidean prompts and keeps all of SAM’s parameters frozen, resulting in a more practical and applicable approach for real clinical scenarios.
Table 4: Performance of SAM with ProMISe (Cross and RN) is tested using GT-base prompts. a Number of Trainable Parameters. b Number of Endoscopy/Dermoscopy images is utilized in the corresponding training set. Some test set images from the corresponding dataset may be seen in the training process.
Benchmarks Kvasir EndoScene ISIC2018
Methods TPa Imagesb mDice mIoU mDice mIoU mDice mIoU
U-Net - 1450/2594 0.818 0.746 0.710 0.627 0.855 0.785
ResUNet++ - 1450/2594 0.821 0.743 0.707 0.624 0.809 0.729
SAM-5P - - 0.750 0.645 0.656 0.582 0.687 0.569
SAM-16P - - 0.719 0.620 0.692 0.613 0.738 0.624
Med2D-5P 184.5M 5838/7935 0.822 0.735 0.722 0.623 0.884 0.813
Med2D-16P 184.5M 5838/7935 0.832 0.748 0.727 0.620 0.893 0.823
Cross-5P 45.6M 1450/2594 0.834 0.744 0.788 0.687 0.742 0.631
Cross-16P 45.6M 1450/2594 0.858 0.777 0.768 0.661 0.819 0.705
RN-5P 23.0M 1450/2594 0.776 0.673 0.705 0.587 0.798 0.688
RN-16P 23.0M 1450/2594 0.846 0.759 0.803 0.735 0.878 0.791
Table 5: Generalization performance of SAM with ProMISe (Cross and RN) is tested using GT-base prompts.
Benchmarks ColonDB ETIS
Methods TPa Imagesb mDice mIoU mDice mIoU
U-Net - 1450/2594 0.512 0.444 0.398 0.335
ResUNet++ - 1450/2594 0.483 0.410 0.401 0.344
SAM-5P - - 0.569 0.482 0.541 0.472
SAM-16P - - 0.548 0.467 0.524 0.455
Med2D-5P 184.5M 5838/7935 0.686 0.576 0.677 0.571
Med2D-16P 184.5M 5838/7935 0.685 0.575 0.622 0.514
Cross-5P 45.6M 1450/2594 0.744 0.636 0.678 0.578
Cross-16P 45.6M 1450/2594 0.732 0.626 0.691 0.591
RN-5P 23.0M 1450/2594 0.664 0.547 0.605 0.508
RN-16P 23.0M 1450/2594 0.735 0.624 0.673 0.566
5.5 Multi-Modality Experiments
To evaluate the multi-modality training potential of our proposed method, we input both endoscopy and dermoscopy together as the training set. Additionally, to verify the stability of our method in multi-modality training, we applied five different random seeds to provide GT-based prompt points during the evaluation. As shown in Fig. 5, the performance of multi-modality training differs only slightly from the results of the corresponding single-modality training (green dashed lines) and is significantly higher than MedSAM-2D (red lines). These experimental results indicate that our IPS method has the potential not only for rapid and low-cost pattern shifting to a single medical domain but also for multiple specific modalities.
Figure 5: Multi-modality training performance of pattern shifting with 16 points (mDice scores). The corresponding results of MedSAM-2D and our proposed IPS are represented by red lines and green dashed lines, respectively, as shown in Table 2 and 3. The blue bars represent the performance of IPS training with both endoscopy and dermoscopy images.
Furthermore, the error bars (standard deviation) indicate that our method has excellent robustness on both Dice and IoU metrics, as the GT-based prompt points generated from five different random seeds produce relatively consistent results. This is particularly meaningful in clinical scenarios, as different clinicians may provide preferred prompt points. Using the method proposed in this paper, they can receive similar results.